OpenStudy (maddie82502):

What is the solution to the inequality? 3x – 11 < –2 A. x > 3 B. x < 3 C. x > –3 D. x < –3

OpenStudy (maddie82502):

@Carpe_Diem77

pooja195 (pooja195):

\[\huge~\rm~3x – 11+11 < –2+11\] -2+11=?

OpenStudy (maddie82502):

13

OpenStudy (anonymous):

looks like pooja is helping you with this one, ill let you continue with poojs

OpenStudy (anonymous):

pooja*

OpenStudy (maddie82502):

k

pooja195 (pooja195):

not quite -2+11 11-2=?

OpenStudy (maddie82502):

9

OpenStudy (texaschic101):

you basically solve it just like equations except for one very important rule....when dividing/multiplying by a negative, the inequality sign changes

pooja195 (pooja195):

good so -2+11=9 \[\huge~\rm~3x < 9] divide both sides by 3 9/3=?

OpenStudy (maddie82502):

3

pooja195 (pooja195):

Good and 3/3=1 so x<9